Akoshia has more money than Komley. If Akoshia gave Komley ¢20, they would have the same amount. While if Komley gave Akoshia ¢22, Akoshia would then have twice as much as Komley. How much does each one actually have?


1
Expert's answer
2021-03-08T19:18:01-0500

Let Akoshia has ¢x, Komley has ¢y, "x>y."

"x-20=y+20"

"x+22=2(y-22)"

Then


"x=y+40"

"y+40+22=2y-44"

"y=106"

"x=146"


Akoshia has ¢146, Komley has ¢106.